OpenStudy (anonymous):

i need a hero !!! What is the sum of a 22–term arithmetic sequence where the first term is 55 and the last term is –29?

OpenStudy (lgbasallote):

too bad @Hero is offline

OpenStudy (anonymous):

just my luck -___-

OpenStudy (lgbasallote):

sorry couldn't resist saying ti

OpenStudy (lgbasallote):

it*

OpenStudy (anonymous):

No hero here, but I can give an answer? We know the average of the series is: \[ \frac{55+(-29)}{2} \]And the number of terms is \(22\), so, we simply state: \[ 22\frac{55+(-29)}{2}=11(55+(-29))=11\cdot 26=286 \]And we are done.

OpenStudy (anonymous):

@lgbasallote You're on a roll, today XD

OpenStudy (anonymous):

whoops ... i meant to ask this quesion .. lol im dumb .. (What is the 15th term of the arithmetic sequence 18, 25, 32, …)

OpenStudy (anonymous):

D: Nooo, what the hell did I do? Oh, all right

jimthompson5910 (jim_thompson5910):

It's an arithmetic sequence, so what's the common difference?

OpenStudy (anonymous):

lol yeahh im sorry, im not completely focused im just trying tto get this done so i dont have to do it tomorrow when im in NY dealing with my grandpas funeral and stuff

jimthompson5910 (jim_thompson5910):

oh i'm so sorry...

jimthompson5910 (jim_thompson5910):

notice how 25 - 18 = 7 and how 32 - 25 = 7 So the common difference is 7 So d = 7

jimthompson5910 (jim_thompson5910):

The first term is 18, so a1 = 18

jimthompson5910 (jim_thompson5910):

You use these two pieces of info to find the nth term an = a1 + (n-1)*d an = 18 + (n-1)*7 an = 18 + 7n - 7 an = 7n - 11 with me so far?

OpenStudy (anonymous):

you completely lost me, but im not trying to learn right now i just need to finish , and its not your fault at all it was his time he was 86

OpenStudy (anonymous):

he lived a good live , a real man among men ♥

OpenStudy (anonymous):

life*